# XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser

### v1

@@ 3.2. Overall Architecture
As shown in Figure 2, we built the XFormParser model based on the multimodal Transformer architecture. The model accepts information from three different modalities: text, position, and vision, which are encoded using text embedding, 2D position embedding and image embedding layers, respectively. Concatenate the text and image embeddings and then add the position embeddings to obtain the input embeddings. The input embedding is encoded by layoutXLM, and the context representation is output through the dense layer. On the one hand, the representation vector is passed through the MLP layer to obtain the entity classification. On the other hand, the entity embedding vector is mined through the Bi-LSTM sequence relationship and then entered into the $MLP_{head}$ or $MLP_{tail}$ according to the type to obtain the entity expression. The score is obtained by Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between two entities.
3.3. layoutXLM and Multimodal Input
@@
layoutXLM adds two new embedding layers, 2-D Position Embedding and Image Embedding, based on BERT. The 2-D Position Embedding utilizes OCR and other technologies to obtain the content and position information of each text block in the document. By considering the top-left corner of the document page as the origin of the coordinate system, the representation of each text block in terms of horizontal coordinate, vertical coordinate, width, and height can be calculated, and the final 2-D Position Embedding is the sum of the representations of the four sub-layers. The Image Embedding divides the image into several blocks based on the bounding boxes of each word in the OCR results and corresponds each block to each word. After normalizing embeddings of multiple modalities, they are input into layoutXLM.
@@
Vectorization of Text Information. Vectorization of text information involves converting textual data into numerical representations that can be processed by machine learning algorithms. This transformation is essential for tasks such as text classification, sentiment analysis, and information retrieval. To achieve this, techniques such as tokenization, word embeddings, and sentence embeddings are employed. Tokenization breaks down the text into smaller units like words or subwords.LayoutXLM uses Roberta Tokenizers (Liu et al., 2019) to map these tokens into dense vectors in a continuous vector space where semantically similar words are positioned closely together. Sentence embeddings, on the other hand, provide a vector representation for entire sentences, capturing the semantic meaning of the sentence as a whole. The final vector representation of the text information is a crucial input for various natural language processing models, enabling them to understand and generate human language more effectively.
@@
Vectorization of Location Information. Two-dimensional position embedding uses OCR and other techniques to obtain the content and position information of each text block in the document. The upper-left corner of the document page is regarded as the origin of coordinates to construct the coordinate system, so that the representation of each text block coordinate corresponding to four embedding sub-layers of coordinates, width and height can be calculated, and the final two-dimensional position embedding is the sum of the embedding of four sub-layers.
@@
Vectorization of Image Feature. In this work, the image embedding is not simply obtained by uniform segmentation, but the text block is used to locate the bounding box of each word in the result, and the image is divided into a number of equal size patches, so as to ensure that each word has the embedding vector of the image containing it.
3.4. Form Cell Classifier.
As shown in Figure 2, the pre-trained model receives input [$[\{b_{1},\ldots,$
$b_{n}\}]$, obtains hidden states $H=[h_{1},\ldots,h_{m}]$, and processes the hidden representation H through a fully connected layer (dense). The hidden representation Hner for the NER task is obtained, and then the classification result $logits_{ner}$ is obtained by a classification MLP. $logits_{ner}$ uses $argmax$ to get the predicted $labels=[L_{1},\ldots,L_{n}]$, $L_{i}$, $L_{i}$ corresponds to the labels of each entity, such as ”QUESTION”, ”ANSWER”, etc. In the input layer, to obtain a better entity representation, Leveraging a more expressive pre-trained model, LayoutXLM introduces coordinate information from bounding boxes during pretraining, for this task, we use label embeddings for each entity. First, we get the pre-trained model output of an entity, denoted as $[h^{re}{begin_{i}},\ldots,h^{re}{end_{i}}]$, the model uses a pooling operation to pool the input representations to obtain the entity representation, and then we concatenate the entity representation and the label embedding $l_{i}$. As input to the document encoder for each semantic entity, the formula is as follows,
(3) $e_{i}=pooling([h_{begin_{i}}^{re},...,h_{end_{i}}^{re}])\oplus l_{i}$
3.5. Entity Relation Classification Decoder.
@@
For the relation discrimination task, in order to make the two tasks share H invisible, there should also be a dense layer for the relation discrimination task, $H_{re}=dense_{re}(H)$. $begin_{i}$ represents the index at which the entity starts in $H^{re}$, $end_{i}$ represents the index at which the entity ends in $H$, the vector from $begin_{i}$ to $end_{i}$ in $H^{re}$ is obtained, and the entity expression in the pre-trained model is obtained by pooling (Reimers and Gurevych, 2019). meanpooling is adopted in this paper, where $(l_{i})$ denotes the vector representation obtained by embedding the entity label $(L_{i})$into the label vector space. In the experiment, it was found that the distribution of entity expression and label embedding was not uniform, which would make it difficult for MLP to learn the importance of both information. Therefore, $e_{i}$ was input into BiLSTM as the encoder to unify, and after the entity passed the encoder, According to the type, it enters the head entity encoder $MLP_{head}$ or the tail entity encoder $MLP_{tail}$, and finally obtains the entity representation. It obtains the score through Biaffine (Nguyen and Verspoor, 2019) to determine whether there is a relationship between the two entities.
3.6. Traning Method
@@
Traning Loss. Since it is a joint model, it needs to be trained together to calculate the loss. Cross entropy is used. The loss formula is as follows:
(4) $Loss_{ner}=\text{CE}(logits_{ner},label_{ner})$
(5) $Loss_{re}=\text{CE}(Score_{relations},label_{re})$
(6) $Loss=Loss_{ner}+Loss_{re}$
@@
where CE is the cross-entropy function. $Loss_{ner}$ is the loss for the entity recognition task, $Loss_{re}$ is the loss for the relationship discrimination task, and the final training Loss is the sum of the two used as the learning target of the joint task.
Warm-up soft label.Soft labeling (Huang et al., 2019), also known as soft targets or probabilistic labeling, is a technique in natural language processing where labels for training data are represented as probability distributions rather than as hard, binary labels. Soft labels provide more information about the uncertainty and distribution of classes, leading to better generalization. In traditional hard labels, each sample can only belong to one category, while using soft labels can represent situations where a sample may belong to multiple categories.
@@
We propose a warm-up soft label based on the warming mechanism. In the beginning stage of training, hard labels are used to supervise the model to converge quickly. In the middle and later stages of training, soft labels are used to help model training. During the transition period between soft and hard labels, a warming mechanism is added. , and continuously increase the weight of soft tags.
This is useful for tasks such as multi-label classification and object detection. For the $logits_{ner}$ obtained in the entity classification task, the soft label is calculated and the soft label embedding is calculated. The specific formula is as follows,
(7) $LE_{sl}=\frac{softmax(logits_{ner})\cdot LE_{weight}}{N}$
where $N$ is the number of tags, $LE_{weight}$ is the weight of the word list that stores tag embeddings, $logits_{ner}$ first obtains the distribution probability through a layer of softmax, and $softmax(logits_{ner})$ Perform dot multiplication with $LE_{weight}$ to get a weighted label embedding vector. Finally, the weighted label embedding vector is divided by the number of samples in the dataset $N$ to obtain the average value of the label embedding vector $LE_{softlabel}$, $LE_{sl}$ for short.
@@
In actual experiments, it was found that the training effect was not good in the first half of model training. Although soft label contains more information, it cannot guide the model to learn tasks in the early stage of training, so hard label is still used in the early stage of model training. After the model has been trained to have preliminary capabilities, then use soft label training, and provide a transition for the conversion of hard label and soft label. In fact, the final label embedding calculation method is as follows:
(8) $\alpha=\min(1,(epoch_{now}-epoch_{start}/epoch_{warm}))$
(9) $LE=\begin{cases}LE_{hl},&epoch_{now}\leq epoch_{start}\\ LE_{hl}\times(1-\alpha)+LE_{sl}\times\alpha,&epoch_{now}>epoch_{start}\\ \end{cases}$
where $\alpha$ is a parameter that decays sublinearly with training. where $epoch_{now}$ represents the current training epoch, $epoch_{start}$ denotes the starting epoch, and $epoch_{warm}$ is a predefined value determining the warm-up duration. The parameter $\alpha$ acts as a scaling factor for the learning rate, ensuring a gradual transition from an initial learning rate to the desired rate, as the training progresses.
Table 1. Training configurations for XFormParser
@@
Evaluation metric F1 score
Table 2. Language-specific fine-tuning F1 accuracy on FUNSD and XFUND (fine-tuning on X, testing on X).“SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
@@
Table 3. Multi-Language fine-tuning accuracy (F1) on the XFUND dataset (fine-tuning on 8 languages all, testing on X), where “SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
@@
4. Experiments and Discussion
4.1. Experiment Settings
Datasets
FUNSD (Jaume et al., 2019) (Guillaume Jaume, 2019) is a scanned document dataset for form parsing. It consists of 149 training samples and 50 test samples with various layouts.XFUND (Xu et al., 2021) (Xu et al., 2022) is a multilingual form parsing benchmark. It includes 1,393 fully annotated forms in 7 languages. Each language contains 199 forms, with 149 forms in the training set and 50 forms in the test set.
XFUND and FUNSD datasets have two tasks: Semantic Entity Recognition, where BIO labeling format is used for sequence labeling of each entity, and Relation Extraction, where all possible pairs of given semantic entities are generated to gradually construct a candidate set of relations, identifying all entity pairs with existing associations.
In this paper, the constructed InDFormBench dataset is analyzed, and compared with the scale of the public dataset. We can see that InDFormBench has the following advantages compared with the other two datasets:
1) Larger data size: The training set of the data set of this paper contains 422 samples, and the validation set and the test set contain 70 samples each. In contrast, the training and validation sets of the FUNSD dataset and the XFUND Chinese dataset both have only 150 samples. Larger data sets can provide more samples for the model to learn and train, which helps improve the generalization ability and performance of the model.
2) Chinese language type: Compared with FUNSD data set which is English language type, the data set of this paper is also Chinese language type. For the research or application scenarios of Chinese document processing, the data set of this paper is more in line with the actual needs, because it better reflects the characteristics and language structure of Chinese documents.
3) Complete data set partition: The training set, validation set and test set of the data set in this paper are clearly divided, containing different numbers of samples respectively. This complete dataset partitioning ensures that the model is trained, validated, and tested on different subsets of the data, thus accurately evaluating the performance and robustness of the model.
Table 4. Cross-lingual zero-shot transfer F1 accuracy on FUNSD and XFUND (fine-tuning on FUNSD, testing on X).
@@ 4.1. Experiment Settings
XFormParser 91.24 74.02 81.77 83.02 79.60 78.61 80.18 81.01 81.18 (+9.83)
4.2. Language-specific Fine-tuning
Table 2 presents the results of specific language fine-tuning experiments on the XFUND and FUNSD public datasets. Each column in the table represents a different language (FUNSD, ZH, JA, ES, FR, IT, DE, PT). The table compares different models, including XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, LayoutLMv3BASE, LiLT[InfoXLM]BASE, GeoLayoutLM, and XFormParserBASE. The numerical values in the table indicate the performance metrics of each model fine-tuned on specific languages. XFormParser demonstrates superior performance compared to other models in most languages. On the FUNSD dataset, XFormParser achieves a performance of 92.46%. XFormParser also exhibits strong performance on the XFUND datasets in different languages, with average performance metrics of 89.04% (SER) and 90.54% (RE).
Table 5. Ablation study of our model using LayoutXLM as the backbone on the FUNSD and XFUND (ZH). The symbol EN denotes FUNSD and ZH means Chinese language.
@@ 4.2. Language-specific Fine-tuning
4.3. Multi-Language Fine-tuning
XFUND dataset inspired Multilingual fine-tuning Task, refers to multilingual task fine-tuning on XFUND dataset, all trained on 8 languages and tested on a specific language, where ”SER” stands for semantic entity recognition and ”RE” stands for relation extraction. As shown in Table 3, in the multilingual training task, XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, and XFormParserBASE all show stronger ability than but language fine-tuning, obtaining higher F1 score. In the case of multiple languages, XFormParserBASE shows a more powerful extraction ability. XFormParser model performs best in the multi-language fine-tuning experiment. The average F1 accuracy of SER task is 91.67%. For RE task, XFormParser model also achieves the best performance in multi-language fine-tuning experiments, with an average F1 of 95.89%. These results underscore the efficacy of the XFormParser model in handling multi-language tasks within the XFUND dataset, particularly in comparison to other baseline and advanced models. With more data, our model demonstrates stronger learning capabilities.
4.4. Zero-shot Fine-tuning
Previous experiments illustrate that our method achieves improvements using full training samples. We explored the transferability of our model structure and found that it will gain strong transferability on RE tasks. Thus, we compare with the previous SOTA model LiLT on few-shot settings. The experimental results in Table 4 indicate that the average performance of XFormParser still outperforms the SOTA model LiLT and GOSE. In the task of SER, XFormParser exhibited the highest F1 scores across all languages, achieving an average of 71.35%, with notable improvements in Chinese (72.00%) and Portuguese (73.46%). In the task of RE, the highest performance in the RE task was again seen with XFormParser, which achieved an impressive average F1 score of 81.18%, showing its robustness in relation extraction across different languages. the XFormParser model consistently outperforms other models in both the SER and RE tasks across various languages, highlighting its superior cross-lingual transfer capabilities, and it can improve the generalization of the model.
4.5. Ablation Study
Table 6. For the ablation experiments of epoch start and epoch warm, epoch start refers to the epoch round when soft label is started, epoch warm refers to whether to use the transition mechanism, and the total training rounds are 100. The first line refers to the method of warm up soft label that is not applicable.
@@ 4.5. Ablation Study
1) w/o multi-task. In this variant, we try to remove the multi-task training method and only use SER or RE for training. This change causes a significant performance decay. The results shown in Table 5 suggest that training two tasks at the same time has the same effect on both SER and RE tasks. With improvement, two tasks that share pre-trained model parameters will learn cross information that improves this task.
@@
2) w/o Decoder. In this variant, we remove the decoder. This change causes a significant performance decay. This suggests the injection of an additional decoder can guide the powerful decoding capabilities of entities and provide strong dependencies for relationship classification.
@@
3) w/o Warm-up soft label. In this variant, we remove the training method Warm-up soft label from XFormParser. This change means the model only uses hard label due to the training. The results shown in Table 5 indicate that Warm-up soft label can improve the effect of the model and prevent the model from overfitting. Experimental data Table 6 indicates that as epoch start increases, model performance initially improves but then decreases. This trend suggests that the model learns sufficient data features by a certain stage, and further training with Warm-up soft label does not enhance performance and may even lead to overfitting.
4.6. Visualization Display

### v2

@@ 3.2 Overall Architecture
As shown in Figure 2, we built the XFormParser model based on the multimodal Transformer architecture. The model accepts information from three different modalities: text, position, and vision, encoded using text embedding, 2D position embedding, and image embedding layers, respectively. Concatenate the text and image embeddings and then add the position embeddings to obtain the input embeddings. The input embedding is encoded by layoutXLM, and the context representation is output through the dense layer. On the one hand, the representation vector is passed through the MLP layer to obtain the entity classification. On the other hand, the entity embedding vector is mined through the Bi-LSTM sequence relation and then entered into the $MLP_{head}$ or $MLP_{tail}$ according to the type to obtain the entity expression. The $Score$ is obtained by Biaffine to determine whether there is a relation between two entities.
3.3 PTM and Multimodal Input
@@
layoutXLM adds two new embedding layers, 2-D Position Embedding and Image Embedding, based on BERT. The 2-D Position Embedding corresponds to text blocks with the content and position information in the document, both content and position information obtained by OCR and other technologies. By considering the top-left corner of the document page as the origin of the coordinate system, the representation of each text block in terms of horizontal coordinate, vertical coordinate, width, and height can be calculated, and the final 2-D Position Embedding is the sum of the representations of the four sub-layers. The image embedding divides the image into several blocks based on the bounding boxes of each word in the OCR results, and each block corresponds to each word. After normalizing embeddings of multiple modalities, they are input into layoutXLM.
@@
Vectorization of Text Information. Tokenization, word embeddings, and sentence embeddings mainly realize the vectorization of text information. Among them, tokenization breaks down text into smaller units such as words or subwords. LayoutXLM uses RoBERTa Tokenizers to map these tokens into dense vectors in a continuous vector space where semantically similar words are positioned closely together. Sentence embeddings are also used, provide a vector representation for entire sentences, capturing the semantic meaning of the sentence as a whole.
@@
Vectorization of Location Information. Two-dimensional position embedding uses OCR and other techniques to obtain the content and position information of each text block in the document. The upper-left corner of the document page is regarded as the origin of constructing the coordinate system so that the embeddings of each text block corresponding to XY coordinates, width, and height can be calculated, and the final two-dimensional position embedding is the sum of the four embeddings.
@@
Vectorization of Image Feature. In this work, the image embedding is not simply obtained by uniform segmentation, but the text block is used to locate the bounding box of each word in the result, and the image is divided into several equal size patches, to ensure that each word has embedding vector of the image containing it.
3.4 Semantic Entity Recognizer
As shown in Figure 2, the PTM receives multi-modal input $B$=$[{b_{1},\ldots,b_{n}}]$, obtains hidden states $H$, and processes the hidden representation $H$ through a fully connected layer (Dense). The hidden representation $H_{ser}$ for the SER task is obtained, and then the classification result $logits_{ser}$ is obtained by a classification MLP. The formulas are expressed as follows:
$H=PLM(B)$ (3)
$H_{ser}=Dense_{ser}(H)$ (4)
$logits_{ser}=MLP_{ser}(H_{ser})$ (5)
Then $logits_{ser}$ uses $argmax$ to get the predicted vectors $P$=$[p_{1},\ldots,p_{n}]$, $p_{i}$ corresponds to the label $l_{i}$ of $i$-th entity, such as "QUESTION", "ANSWER", etc.
3.5 Relation Extraction Decoder
For the RE task, we first get the PTM output of the $i$-th entity, denoted as $H_{re}$. The model then uses a pooling operation Reimers and Gurevych (2019) on $H_{re}$ to obtain the embeddings of the entities. Then we concatenate the entity embeddings and the corresponding label embeddings $p_{i}$ predicted by the SER. As the output of the encoder for each semantic entity, the formula is as follows:
$e_{i}=pooling(H_{re})\oplus p_{i}$ (6)
Note that for two tasks to implicitly share $H$, $H_{re}$ needs to be obtained using a Dense layer identical to Eq. (4), that is, $H_{re}$= $Dense_{re}(H)$. In addition, mean-pooling is experimentally verified to be the most effective pooling operation applied in Eq. (6).
@@ 3.5 Relation Extraction Decoder
In the experiment, it was found that the distribution of entity expression and label embedding was not uniform, which would make it difficult for MLP to learn the importance of both information. Therefore, $e_{i}$ was input into the Bi-LSTM decoder to unify, and after the entity passed the decoder, according to the type, it entered the head entity decoder $MLP_{head}$ or the tail entity decoder $MLP_{tail}$, and finally produces the entity representation. It obtains the score through Biaffine to determine whether there is a relation between the two entities.
3.6 Traning Method
@@
Training Loss. Since it is a joint model, the team must be trained to calculate the loss. The loss formula is as follows:
$Loss_{ser}=\text{CE}(logits_{ser},L_{ser})$ (7)
$Loss_{re}=\text{CE}(Score_{Biaffine},L_{re})$ (8)
$Loss=Loss_{ser}+Loss_{re}$ (9)
@@
where CE is the cross-entropy function. $Loss_{ser}$ is the loss for the SER task, $Loss_{re}$ is the loss for the RE task, and the final training Loss is the sum of the two used as the learning target of the joint task.
@@
Warm-up Soft Label. We propose an improved warm-up soft label Huang et al. (2019) based on the warming mechanism, which can effectively improve the performance when applied to fine-tuning form parsing models. In the beginning stage of training, hard labels are used to supervise the model so that it can converge quickly. In the middle and later stages of training, soft labels are used to help model training. During the transition period between soft and hard labels, a warming mechanism is added, and the weight of soft tags is continuously increased. This is useful for tasks such as multi-label classification.
@@
In actual experiments, it was found that the training effect was not good in the first half of model training. Although soft labels contain more information, they cannot guide the model in learning tasks in the early stage of training, so hard labels are still used in the early stage of model training. After the model has been trained to have preliminary capabilities, then use soft label training, and provide a transition for the conversion of hard label and soft label. For the construction method of the warm-up soft label and related training parameter Settings, please refer to Appendix A.
@@
Evaluation metric ref. B.4
Table 1: Training configurations for XFormParser
@@
Table 2: Language-specific fine-tuning F1 accuracy on FUNSD and XFUND (fine-tuning on X, testing on X).“SER” denotes the semantic entity recognition, and “RE” denotes the relation extraction.
@@
Table 3: Multi-language fine-tuning accuracy (F1) on the XFUND dataset (fine-tuning on 8 languages all, testing on X), where “SER” denotes the semantic entity recognition and “RE” denotes the relation extraction.
4 Experiments and Discussion
4.1 Experiment Settings
Datasets. FUNSD Guillaume Jaume (2019) is a scanned document dataset for form parsing. It is a subset of the RVL-CDIP Harley et al. (2015) and consists of 149 training samples and 50 test samples with various layouts. XFUND Xu et al. (2022) is a multilingual form parsing benchmark. It includes 1,393 fully annotated forms in 7 languages. Each language contains 199 forms, with 149 forms in the training set and 50 forms in the test set.
XFUND and FUNSD datasets have two tasks: SER, where BIO labeling format is used for sequence labeling of each entity, and RE, where all possible pairs of given semantic entities are generated to gradually construct a candidate set of relations, identifying all entity pairs with existing associations.
In this paper, InDFormSFT is constructed based on the Chinese-English context and eight application scenarios, the training set contains 422 samples, and the validation set and test set contain 70 samples each. Compared with FUNSD and XFUND datasets, larger datasets can provide more samples for the model to learn and train, which helps to improve the generalization ability and performance of the model. Please refer to Appendix B for the construction process of InDFormSFT.
@@ 4.1 Experiment Settings
XFormParser 91.24 74.02 81.77 83.02 79.60 78.61 80.18 81.01 81.18 (+13.89)
Table 4: Cross-lingual zero-shot transfer F1 accuracy on FUNSD and XFUND (fine-tuning on FUNSD, testing on XFUND).
4.2 Language-specific Fine-tuning
Table 2 presents the results of specific language fine-tuning experiments on the XFUND and FUNSD public datasets. Each column in the table represents a different language (FUNSD, ZH, JA, ES, FR, IT, DE, PT) Xu et al. (2022). The table compares different models, including XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, LayoutLMv3BASE, LiLT[InfoXLM]BASE, GeoLayoutLM, and XFormParserBASE. The numerical values in the table indicate the performance metrics of each model fine-tuned in specific languages. XFormParser demonstrates superior performance compared to other models in most languages. On the FUNSD dataset, XFormParser achieves a performance of 92.46%. XFormParser also exhibits strong performance on the XFUND datasets in different languages, with average performance metrics of 89.04% (SER) and 90.54% (RE).
@@ 4.2 Language-specific Fine-tuning
Table 5: Ablation study of our model using LayoutXLM as the backbone, InDFormSFT as training dataset and test on the FUNSD and XFUND (ZH). The symbol EN denotes FUNSD and ZH means Chinese language.
4.3 Multi-language Fine-tuning
XFUND dataset inspired Multilingual fine-tuning Task, refers to multilingual task fine-tuning on XFUND dataset, all trained on 8 languages and tested on a specific language. As shown in Table 3, in the multilingual training task, XLM-RoBERTaBASE, InfoXLMBASE, LayoutXLMBASE, and XFormParserBASE all show stronger ability than language-specific fine-tuning, obtaining higher F1 score. In the case of multiple languages, XFormParserBASE shows a more powerful extraction ability. XFormParser performs best in the multi-language fine-tuning experiment. The average F1 accuracy of SER task is 91.67%. For RE task, XFormParser also achieves the best performance in multi-language fine-tuning experiments, with an average F1 of 95.89%. These results underscore the efficacy of the XFormParser in handling multi-language tasks within the XFUND dataset, particularly in comparison to other baseline and advanced models. With more data, our model demonstrates stronger learning capabilities.
4.4 Zero-shot Fine-tuning
Previous experiments illustrate that our method achieves improvements using full training samples. We explored the transferability of our model structure and found that it will gain strong transferability on RE tasks. Thus, we compare with the previous SOTA model LiLT on few-shot settings. The experimental results in Table 4 indicate that the average performance of XFormParser still outperforms the SOTA model LiLT and GOSE. In the task of SER, XFormParser exhibited the highest F1 scores across all languages, achieving an average of 71.35%, with notable improvements in Chinese (72.00%) and Portuguese (73.46%). In the task of RE, the highest performance in the RE task was again seen with XFormParser, which achieved an impressive average F1 score of 81.18%, showing its robustness in relation extraction across different languages. the XFormParser consistently outperforms other models in both the SER and RE tasks across various languages, highlighting its superior cross-lingual transfer capabilities, and it can improve the generalization of the model.
4.5 Ablation Study
@@
Table 6: For the ablation experiments of epoch start and epoch warm, epoch start refers to the epoch round when the soft label is started, epoch warm refers to whether to use the transition mechanism, and the total training rounds are 100. The first line refers to the method of warm-up soft label that is not applicable.
@@
1) w/o multi-task. In this variant, we try to remove the multi-task training method and only use SER or RE for training. This change causes a significant performance decay. The results shown in Table 5 suggest that training two tasks at the same time has the same effect on both SER and RE tasks. With improvement, two tasks that share PTM parameters will learn cross-information that improves this task.
@@
2) w/o Decoder. In this variant, we remove the decoder. This change causes a significant performance decay. This suggests the injection of an additional decoder can guide the powerful decoding capabilities of entities and provide strong dependencies for relation classification.
@@
3) w/o Warm-up soft label. In this variant, we remove the training method Warm-up soft label from XFormParser. This change means the model only uses hard labels due to the training. The results shown in Table 5 indicate that a Warm-up soft label can improve the effect of the model and prevent the model from overfitting. Experimental data Table 6 indicates that as epoch starts increases, model performance initially improves but then decreases. This trend suggests that the model learns sufficient data features by a certain stage, and further training with a Warm-up soft label does not enhance performance and may even lead to overfitting.
4.6 Visualization Display